# RPCBF: Constructing Safety Filters Robust to Model Error and Disturbances via Policy Control Barrier Functions

### v2

@@ Abstract
Control Barrier Functions (CBFs) have proven to be an effective tool for performing safe control synthesis for nonlinear systems. However, guaranteeing safety in the presence of disturbances and input constraints for high relative degree systems is a difficult problem. In this work, we propose the Robust Policy CBF (RPCBF), a practical method of constructing CBF approximations that is easy to implement and robust to disturbances via the estimation of a value function. We demonstrate the effectiveness of our method in simulation on a variety of high relative degree input-constrained systems. Finally, we demonstrate the benefits of RPCBF in compensating for model errors on a hardware quadcopter platform by treating the model errors as disturbances.
@@ I Introduction and Related Works
In the realm of autonomous systems, providing safety guarantees is crucial, especially in critical applications such as autonomous driving [1] and healthcare robotics [2]. Control Barrier Functions (CBFs) [3, 4] have proven to be an effective tool to maintain and certify the safety of dynamical systems. In particular, they can be applied as a Safety Filter (SF) that minimally modifies arbitrary control inputs to ensure safety, making them especially valuable when integrated with learning-based controllers.
@@
Despite their theoretical advantages, significant challenges remain in the practical application and construction of CBFs. First, constructing CBFs is non-trivial, specifically for high relative degree systems with input constraints. Second, the safety guarantees of  CBF-based controllers depend on having an accurate system model, which is rarely the case for systems in real life. This can result in the safety of these controllers being sensitive to model uncertainties.
Learning Control Barrier Functions. To minimize reliance on extensive domain knowledge, a recent trend is to learn neural CBFs that approximate CBFs using Neural Networks (NNs) [5, 6, 7, 8, 9, 10, 11, 12, 13]. Leveraging the flexibility of NNs, neural CBFs have been successfully applied to high-dimensional systems, including multi-agent control scenarios [14, 11]. Furthermore, they have been extended to handle parametric uncertainties [15] and obstacles with unknown dynamics [16]. Although utilizing NNs as CBFs offers universal approximation capabilities, it requires their certification as CBF to provide safety guarantees and limits their interpretability.
Furthermore, using a naive approach of learning neural CBFs by minimizing a loss that encourages the CBF conditions can lead to a small or even empty forward-invariant set. Thus, [8] present a method to construct CBFs using policy evaluation of any policy. They show that the policy value function is a CBF and learn an NN approximation of it. In this setting, the policy value function represents the maximum-over-time constraint violation and is generally a measure of how good a particular state is for a system following a specific policy over infinite time. However, their approach does not consider uncertainties in the system dynamics.
@@
Figure 1: We propose the Robust Policy Control Barrier Function (RPCBF), which approximates the value function $V^{h,\pi}$ of a system under bounded disturbances for policy $\pi$ online. The zero sublevel set of the RPCBF is a robust controlled-invariant set (CI) and can be used online as a safety filter to ensure safety for any unsafe nominal policy. We demonstrate its superior performance compared to a safety filter using a PCBF on a quadcopter platform with model errors treated as disturbances.
@@
Robust Safety. Controllers robust to disturbances in the real world are important for safe autonomous systems. This has been studied before in robust CBFs [17, 15, 18] which guarantee safety under bounded disturbances. However, constructing robust CBFs is more difficult than regular CBFs, especially under input constraints. Hamilton-Jacobi reachability analysis [19] can be used to compute robust control-invariant sets, which can then be subsequently used for constructing robust CBFs [20, 21, 22, 23]. However, reachability analysis in itself is challenging, with grid-based PDE solvers being limited to state dimensions less than $5$ [24], while deep learning-based solvers [25, 26, 27, 28] require subsequent neural-network verification to check for solution accuracy.
To address these challenges, we propose a practical approach to construct robust CBFs that guarantee safety under worse-case bounded disturbances through policy evaluation of any policy, by estimating the policy value function, which has been demonstrated to be a valid CBF in [8]. Our approach leverages finite horizon policy rollouts and extends the method to handle the robust case.
@@
    We propose a method of constructing robust CBFs using the robust policy value function and subsequently propose a real-time approximation that can be used online.
@@
    We examine the benefits of robust CBFs on a hardware quadcopter, where the additional robustness to model error is crucial to prevent collisions.
@@ II-A Problem Statement
$\dot{\mathbf{x}}=f(\mathbf{x},\mathbf{d})+g(\mathbf{x},\mathbf{d})\mathbf{u},$ (1)
with state $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$, control input $\mathbf{u}\in\mathcal{U}\subseteq\mathbb{R}^{m}$, (unknown) bounded smooth disturbance $\mathbf{d}_{\mathrm{min}}\leq\mathbf{d}\leq\mathbf{d}_{\mathrm{max}}$ with $\mathbf{d}_{\mathrm{min}},\mathbf{d}_{\mathrm{max}}\in\mathbb{R}^{d}$ and $f$, $g$ locally Lipschitz continuous. Let $\mathcal{A}\subset\mathcal{X}$ denote the set of states to be avoided. In this paper, we address the following SF synthesis problem:
@@
Given a dynamical system (1) and an avoid set $\mathcal{A}\subset\mathcal{X}$, find a control policy $\pi_{\mathrm{filt}}:\mathcal{X}\rightarrow\mathcal{U}$ that ensures the system state remains outside the avoid set $\mathcal{A}$ while staying close to a performant but possibly unsafe nominal policy $\pi_{\mathrm{nom}}:\mathcal{X}\rightarrow\mathcal{U}$:
$\min_{\pi}$ $\lVert\pi_{\mathrm{filt}}-\pi_{\mathrm{nom}}\rVert$ (2)
[…]
@@ II-B Safety Filters using Control Barrier Functions
$\dot{\mathbf{x}}=f(\mathbf{x},0)+g(\mathbf{x},0)\mathbf{u}.$ (3)
Let $B:\mathcal{X}\rightarrow\mathbb{R}$ be a continuously differentiable function, with $\mathcal{C}=\{\mathbf{x}\in\mathcal{X}\,|\,B(\mathbf{x})\leq 0\}$ as its 00-sublevel set. Let $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ be an extended class-$\kappa_{\infty}$ function [^1] Extended class-$\kappa_{\infty}$ is the set of continuous, strictly increasing functions $\alpha:(-\infty,\infty)\rightarrow(-\infty,\infty)$ with $\alpha(0)=0$.. Then, $B$ is a CBF for the undisturbed system (3) on $\mathcal{X}$ [3] if
@@
renders $\mathcal{C}$ forward-invariant [3]. In other words, there exists an $\mathbf{u}\in\mathcal{U}$ such that any trajectory starting within $\mathcal{C}$ remains in $\mathcal{C}$. Asymptotic stability of $\mathcal{C}$ can be achieved by extending (4b) to hold for all $\mathbf{x}\in\mathcal{X}$ [29]. Since the right hand side of condition (4b) is linear in $\mathbf{u}$, given a CBF $B$ where (4b) is satisfied, we can solve 1 for (3) using the following Quadratic Program (QP)-based controller:
@@
[…]
While CBFs can be applied to guarantee safety for a known undisturbed system, two major challenges remain:
@@
    How to synthesize a $B$ that satisfies (4b) so it is a CBF?
@@
    How do we guarantee that using the CBF for control synthesis results in safe controllers when applied to the original disturbed system?
III Robust Policy Control Barrier Functions
To address the above two challenges, we leverage the insight from [8] that CBFs can be constructed by deriving the policy value function through the evaluation of any policy. Rather than approximating the policy value function with an NN as in [8], we propose an alternative that avoids NNs by instead performing a finite-horizon numerical approximation. We further extend this approach to the robust case and introduce Robust Policy CBFs (RPCBFs). We begin by revisiting the formulation of PCBFs and describe our extensions.
@@ III-A Constructing CBFs via Policy Evaluation
We denote by $\mathbf{x}_{t}^{\pi}$ the resulting state at time $t$ when starting from the initial state $\mathbf{x}_{0}$ and following policy $\pi:\mathcal{X}\rightarrow\mathcal{U}$. Furthermore, we define the maximum-over-time value function for the undisturbed system in (3) as
@@
As stated in [8, Theorem 1], the policy value function $V^{h,\pi}_{\infty}$ is a CBF for the undisturbed system in (3) for any $\pi$, since $V^{h,\pi}_{\infty}$ satisfies the following two inequalities $\forall\mathbf{x}\in\mathcal{X}$
@@
which imply (4a) and (4b). For details, we refer to [8]. The key intuition here is that $V^{h,\pi}_{\infty}$ provides an upper bound on the worst future constraint violation $h$ under the optimal policy since the optimal policy will do no worse than $\pi$. Thus, CBFs can be constructed via policy evaluation of any policy.
@@ III-B Finite Horizon Approximation of PCBFs
A key challenge with the value function $V^{h,\pi}_{\infty}$ is that its definition requires an infinite-horizon. While [8] tackles this problem by using an NN to learn $V^{h,\pi}_{\infty}$ with a loss derived using dynamic programming, we take a different approach and perform a finite-horizon approximation that can be computed without the use of NN, enabling a more in-depth analysis of the resulting safety guarantees. Expanding $V^{h,\pi}_{\infty}$:
@@
$\approx\underbrace{\sup_{0\leq t<T}h(\mathbf{x}_{t}^{\pi})}_{\coloneqq V^{h,\pi}_{T}(\mathbf{x}_{0})},$ (10)
where the approximation is made by dropping the $V_{\infty}^{h,\pi}(\mathbf{x}_{T})$ “tail”. The question is then whether the finite-horizon approximation $V_{T}^{h,\pi}$ is a CBF and has safety guarantees.
We can at least answer this in the affirmative when the approximation in (10) is an equality, i.e., the maximum occurs in $[0,T]$. We state this in the following theorem.
@@
Suppose there exists a $\bar{T}<\inf$ such that
$\operatorname*{arg\,max}_{t\geq 0}h(\mathbf{x}_{t}^{\pi})<\bar{T},\quad\forall\mathbf{x}_{0}\text{ where }V_{T}^{h,\pi}(\mathbf{x}_{0})\leq 0$ (12)
Then, $V^{h,\pi}_{T}$ is a CBF for $T\geq\bar{T}$.
@@
(12) implies (11) for $T\geq\bar{T}$. The proof then follows from Theorem 1. ∎
@@
The finite-horizon approximation here is closely related to the use of MPC by practitioners. More precisely, although a terminal constraint set is often required to theoretically guarantee recursive feasibility of finite-horizon MPC [30, 31], practitioners often apply MPC without the use of such a terminal constraint set to wide success [32, 33, 34, 35]. Our decision to drop the $V^{h,\pi}_{\infty}(\mathbf{x}_{T}^{\pi})$ term can be viewed as being similar to dropping the terminal constraint set. Another similarity is the choice of horizon $T$. Namely, recursive feasibility holds in MPC given a sufficiently large horizon [36], similar to corollary 1.
Remark (Connections to Backup Controller / CBFs).
Since the zero sublevel set of $V^{h,\pi}_{\infty}$ is a controlled-invariant set under $\pi$, (9) can also be seen as Backup CBF [37, 38] with backup controller $\pi$. Unlike this (and other similar approaches [39]), our approach replaces the need for a known forward-invariant set with the requirement of a sufficiently long horizon $T$.

### v3

# Safety on the Fly: Constructing Robust Safety Filters via Policy Control Barrier Functions at Runtime

@@ Abstract
Control Barrier Functions (CBFs) have proven to be an effective tool for performing safe control synthesis for nonlinear systems. However, guaranteeing safety in the presence of disturbances and input constraints for high relative degree systems is a difficult problem. In this work, we propose the Robust Policy CBF (RPCBF), a practical approach for constructing robust CBF approximations online via the estimation of a value function. We establish conditions under which the approximation qualifies as a valid CBF and demonstrate the effectiveness of the RPCBF-safety filter in simulation on a variety of high relative degree input-constrained systems. Finally, we demonstrate the benefits of our method in compensating for model errors on a hardware quadcopter platform by treating the model errors as disturbances. Website including code: www.oswinso.xyz/rpcbf/
@@ I Introduction and Related Works
In the realm of autonomous systems, providing safety guarantees is crucial, especially in critical applications such as autonomous driving and healthcare robotics. Control Barrier Functions (CBFs) [6, 45] have proven to be an effective tool to maintain and certify the safety of dynamical systems. In particular, they can be applied as a Safety Filter (SF) that minimally modifies arbitrary control inputs to ensure safety, making them especially valuable when integrated with learning-based controllers.
@@
Despite their theoretical advantages, significant challenges remain in the practical application and construction of CBFs. First, constructing CBFs is non-trivial, specifically for high relative degree systems with input constraints. Second, the safety guarantees of  CBF-based controllers depend on having an accurate system model, which is rarely the case for systems in real life. This makes the safety guarantees of such controllers sensitive to model uncertainties.
@@
Figure 1: We propose the Robust Policy Control Barrier Function (RPCBF), which approximates the robust value function $V^{h,\pi}$ of a system under bounded disturbances for policy $\pi$ at runtime. The zero sublevel set of $V^{h,\pi}$ is a robust controlled-invariant (CI) set. We apply a RPCBF-Safety Filter (SF) to ensure safety for any unsafe nominal policy, demonstrating superior performance over the non-robust PCBF-SF on a quadcopter with model errors treated as disturbances.
Learning Control Barrier Functions. To minimize reliance on extensive domain knowledge, a recent trend is to learn neural CBFs that approximate CBFs using Neural Networks (NNs) [13, 23, 35, 32, 52, 37, 43]. Neural CBFs have been successfully applied to high-dimensional systems, including multi-agent control scenarios [52, 31], and have been extended to handle parametric uncertainties [14] and obstacles with unknown dynamics [51]. Although using NNs as CBFs offers universal approximation capabilities, it requires certifying them as valid CBFs to ensure safety guarantees and limits their interpretability. Furthermore, using a naive approach to learning neural CBFs by minimizing a loss that encourages the CBF conditions can lead to a small or even empty forward-invariant set. Thus, [35] presents a method to construct CBFs using policy evaluation of any policy. They show that the policy value function is a CBF and learn an NN approximation. In this setting, the policy value function represents the maximum-over-time constraint violation, indicating how suitable a state is for a system following a specific policy. However, their approach does not consider uncertainties in the system dynamics.
@@
Robust Safety. Controllers that are robust to disturbances are essential for ensuring the safety of autonomous systems in the real world. This has been studied before in robust CBFs [14, 18, 12], which guarantee safety under bounded disturbances. However, constructing robust CBFs is inherently more difficult than constructing standard CBFs, especially under input constraints. Hamilton-Jacobi reachability analysis [26] can be used to compute robust control-invariant sets, which can then be subsequently used for constructing robust CBFs [11, 41, 39, 40]. However, reachability analysis in itself is challenging, with grid-based partial Differential Equation (DE) solvers being limited to state dimensions below five [27], while deep learning-based solvers [7, 17, 34] require subsequent NN verification to check for solution accuracy. Moreover, both learning-based CBF approaches and deep learning-based reachability solvers depend on predefined system dynamics and disturbance assumptions, which are difficult to adapt, limiting their flexibility once deployed, as retraining cannot be performed on the system. Thus, [22] train a value network with the avoidance set and disturbance bounds as inputs, however this increases training data requirements and complicates evaluating how well the learned network represents the true value function.
As an alternative to robust safety, other works focus on risk-aware safety, which aims to ensure safety with high probability by modeling disturbances probabilistically and incorporating risk measures [2, 24]. Unlike robust methods, which ensure constraint satisfaction for all disturbances within a bounded set but may be overly conservative, risk-aware approaches typically rely on knowledge of the disturbance’s probability distribution. In this work, we focus on robust safety.
We propose a practical approach for constructing a CBF approximation at runtime, which can be derived for any system dynamics and disturbance bounds without requiring (re)training. We establish conditions under which the resulting CBF approximation qualifies as a valid CBF. Our method constructs CBFs by evaluating the value function of any policy, which has been shown to be a valid CBF in [35]. By leveraging finite-horizon policy rollouts, we enable a more detailed analysis of safety guarantees than NN approximations. We apply this approach to construct approximations of robust CBFs.
@@
    We propose a method of constructing (robust) CBFs using the (robust) policy value function and a real-time approximation that can be used at runtime.
@@
    We demonstrate real-time performance and the benefits of our robust CBFs on a hardware quadcopter, where robustness to model errors is key for collision prevention.
@@ II-A Problem Statement
$\dot{\mathbf{x}}_{t}=f(\mathbf{x}_{t},\mathbf{d}_{t})+g(\mathbf{x}_{t},\mathbf{d}_{t})\mathbf{u}_{t},$ (1)
with state $\mathbf{x}_{t}\in\mathcal{X}\subseteq\mathbb{R}^{n}$, control input $\mathbf{u}_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ and unknown, bounded, smooth disturbance $\mathbf{d}_{\mathrm{min}}\leq\mathbf{d}_{t}\leq\mathbf{d}_{\mathrm{max}}$ with $\mathbf{d}_{\mathrm{min}},\mathbf{d}_{\mathrm{max}}\in\mathbb{R}^{d}$ (e.g., estimated from empirical data), where $\mathbf{d}_{t}$ can be time-varying. The functions $f$ and $g$ are assumed to be locally Lipschitz continuous. Let $\mathcal{A}\subset\mathcal{X}$ denote the set of states to be avoided. This paper addresses the following SF synthesis problem:
@@
Given a system (1) and an avoid set $\mathcal{A}\subset\mathcal{X}$, find a control policy $\pi_{\mathrm{filt}}:\mathcal{X}\rightarrow\mathcal{U}$ that keeps the system state outside $\mathcal{A}$ while staying close to a performant, possibly unsafe nominal policy $\pi_{\mathrm{nom}}:\mathcal{X}\rightarrow\mathcal{U}$:
$\min_{\pi_{\mathrm{filt}}}$ $\lVert\pi_{\mathrm{filt}}-\pi_{\mathrm{nom}}\rVert$ (2)
s.t. ${\color[rgb]{0,0,0}\dot{\mathbf{x}}_{t}=f(\mathbf{x}_{t},\mathbf{d}_{t})+g(\mathbf{x}_{t},\mathbf{d}_{t})\pi_{\mathrm{filt}}(\mathbf{x}_{t})}$
$\mathbf{x}_{t}\notin\mathcal{A},\ \forall t\geq 0,$
where $\lVert\cdot\rVert$ is some distance metric.
@@ II-B Safety Filters using Control Barrier Functions
$\dot{\mathbf{x}}_{t}=f(\mathbf{x}_{t},0)+g(\mathbf{x}_{t},0)\mathbf{u}_{t}.$ (3)
Let $B:\mathcal{X}\rightarrow\mathbb{R}$ be a continuously differentiable function, with $\mathcal{C}=\{\mathbf{x}\in\mathcal{X}\,|\,B(\mathbf{x})\leq 0\}$ as its 0-sublevel set. Let $\alpha:\mathbb{R}\rightarrow\mathbb{R}$ be an extended class-$\kappa_{\infty}$ function [^1] Extended class-$\kappa_{\infty}$ is the set of continuous, strictly increasing functions $\alpha:(-\infty,\infty)\rightarrow(-\infty,\infty)$ with $\alpha(0)=0$.. Then, $B$ is a CBF for the undisturbed system (3) on $\mathcal{X}$ [6] if
@@
renders $\mathcal{C}$ forward-invariant [6]. In other words, there exists an $\mathbf{u}\in\mathcal{U}$ such that any trajectory starting within $\mathcal{C}$ remains in $\mathcal{C}$. Asymptotic stability of $\mathcal{C}$ can be achieved by extending (4b) to hold for all $\mathbf{x}\in\mathcal{X}$ [47]. Since the right hand side of (4b) is linear in $\mathbf{u}$, given a CBF $B$, we can solve 1 for (3) using the following Quadratic Program (QP)-based controller:
@@
$\text{s.t.}\ L_{f}B(\mathbf{x})+L_{g}B(\mathbf{x})\mathbf{u}\leq-\alpha(B(\mathbf{x})).$
While CBFs can be applied to guarantee safety for a known undisturbed system, three major challenges remain:
@@
    How do we synthesize a valid CBF that satisfies (4b) for high relative degree systems with input constraints?
@@
    How do we synthesize a robust CBF that ensures safe control for the disturbed system?
  3. 3.
    How can we efficiently derive a CBF at runtime for different system dynamics and disturbance assumptions?
III (Robust) Policy Control Barrier Functions
To address the above challenges, we leverage the insight from [35] that CBFs can be constructed by deriving the policy value function through the evaluation of any policy. Rather than approximating the policy value function with an NN as in [35], we propose a real-time approximation that avoids NNs and can be derived at runtime through a finite-horizon numerical approximation. We further extend this approach to the robust case and introduce Robust Policy Control Barrier Functions (RPCBFs) and subsequently propose a sampling-based approximation that can be derived at runtime. Next, we revisit the formulation of PCBFs and describe our extensions and approximations.
@@ III-A Constructing CBFs via Policy Evaluation
Note that $h(\mathbf{x})>0$ for states that are already in the failure set, whereas $B(\mathbf{x})>0$ for states from which failure is inevitable in the future under the given dynamics and input constraints. In the absence of input constraints, $h$ and $B$ may coincide. We denote by $\mathbf{x}_{t}^{\pi}$ the resulting state at time $t$ when starting from the initial state $\mathbf{x}_{0}$ and following policy $\pi:\mathcal{X}\rightarrow\mathcal{U}$. Furthermore, we define the maximum-over-time value function for the undisturbed system in (3) as
@@
As stated in [35, Theorem 1], the policy value function $V^{h,\pi}_{\infty}$ is a CBF for the undisturbed system in (3) for any $\pi$, since $V^{h,\pi}_{\infty}$ satisfies the following two inequalities $\forall\mathbf{x}\in\mathcal{X}$
@@
which imply (4a) and (4b). For details, we refer to [35]. The key intuition here is that $V^{h,\pi}_{\infty}$ provides an upper bound on the worst future constraint violation $h$ under the optimal policy since the optimal policy will do no worse than $\pi$. Thus, CBFs can be constructed via policy evaluation of any policy. We refer to $\pi$ as the design policy, noting that the nominal policy $\pi_{\mathrm{nom}}$ differs from the design policy.
@@ III-B Finite Horizon Approximation of PCBFs
A key challenge with the policy value function $V^{h,\pi}_{\infty}$ is that its definition requires an infinite-horizon. While [35] tackles this problem by using an NN to learn $V^{h,\pi}_{\infty}$ with a loss derived using dynamic programming, we take a different approach and perform a finite-horizon approximation that can be computed without the use of an NN, enabling a more in-depth analysis of the resulting safety guarantees. Expanding $V^{h,\pi}_{\infty}$:
@@
$\approx\sup_{0\leq t<T}h(\mathbf{x}_{t}^{\pi}){\color[rgb]{0,0,0}\coloneqq V^{h,\pi}_{T}(\mathbf{x}_{0})},$ (10)
where the approximation is made by dropping the $V_{\infty}^{h,\pi}(\mathbf{x}_{T})$ “tail”. The question is then whether the finite-horizon approximation $V_{T}^{h,\pi}$ is a CBF and can provide safety guarantees.
We can at least answer this in the affirmative when the approximation in (10) is an equality, i.e., the maximum occurs in $0,T{\color[rgb]{0,0,0})}$. We state this formally in the following theorem.
@@
Suppose there exists a $\tilde{T}<\inf$ such that
$\operatorname*{arg\,max}_{t\geq 0}h(\mathbf{x}_{t}^{\pi})<\tilde{T},\quad\forall\mathbf{x}_{0}\text{ where }V_{T}^{h,\pi}(\mathbf{x}_{0})\leq 0.$ (12)
Then, $V^{h,\pi}_{T}$ is a CBF for $T\geq\tilde{T}$.
@@
(12) implies (11) for $T\geq\tilde{T}$. Proof from theorem 1. ∎
The value of $\tilde{T}$ depends on the chosen design policy $\pi$. Any policy can be selected, but a conservative Controlled-Invariant (CI) set may result. If $\pi$ is chosen as a controller that steers the system towards a safe, steady-state or CI set within a finite horizon $\tilde{T}$, Corollary 1 holds, thus $V_{T}^{h,\pi}$ is a valid CBF.
If $V^{h,\pi}_{T}(\mathbf{x}_{0})=0$, then applying the design policy $\pi$ exactly guarantees that the system will remain safe for at least the time horizon $0\leq t<T$, even if $V^{h,\pi}_{T}(\mathbf{x}_{0})<V^{h,\pi}_{\infty}(\mathbf{x}_{0})$.
Remark (Connections to Backup Controller / CBFs).
Since the zero sublevel set of $V^{h,\pi}_{\infty}$ is a CI set under $\pi$, (9) can also be seen as Backup CBF [33, 10] with backup controller $\pi$ and no known CI terminal set. Unlike this (and other similar approaches [1]), our approach replaces the need for a known CI set with the requirement of a sufficiently long horizon $T$. Thus, the design policy $\pi$ can be chosen arbitrarily and is not required to steer the system into a CI set. Furthermore, we demonstrate that the naive approximation of $h$ over a time-discretized state trajectory introduces gradient errors. To address this, we present an improved time-discretization using cubic splines in section III-C.
@@
The finite-horizon approximation here is closely related to the use of MPC by practitioners. More precisely, although a terminal constraint set is often required to theoretically guarantee recursive feasibility of finite-horizon MPC [25, 9], practitioners often apply MPC without the use of such a terminal constraint set to wide success [19, 4, 44, 36]. Our decision to drop the $V^{h,\pi}_{\infty}(\mathbf{x}_{T}^{\pi})$ term can be viewed as being similar to dropping the terminal constraint set. Another similarity is the choice of horizon $T$. Namely, recursive feasibility holds in MPC given a sufficiently large horizon [8], similar to Corollary 1. However, the MPC horizon length is limited, as it requires solving a potentially nonlinear and non-convex optimization problem online, with computational complexity typically scaling cubically with $T$ [20]. A key advantage of PCBF-SFs is that they only solve the simpler (CBF-QP), whose computation time is unaffected by $T$, see Section IV-D.
Figure 2: Value Function Gradient Error for Discrete-Time Double Integrator. We highlight the discretized trajectory (top) and corresponding gradient error (bottom) for three different choices of $\Delta t$ (yellow, purple, red). The gradient of the naive discrete-time approximation has large errors and varies with the choice of $\Delta t$. Taking the maximum of the cubic spline leads to much smaller errors.
Remark (Connections to Predictive Safety Filter (PSF) [42]).
The finite-horizon approximation is closely related to the PSF, which implicitly represents the safe set via a finite-horizon MPC problem with terminal constraints or long horizons for recursive feasibility. Unlike PCBF-SFs, the PSF requires solving a potentially nonlinear and nonconvex optimization problem online with complexity scaling cubically in $T$ [20]. PCBF-SFs only require solving the simpler (CBF-QP). However, while the PSF may find a locally optimal solution, the conservativeness of the PCBF-SF depends on $\pi$.